# Learning Human-Aware Robot Policies for Adaptive Assistance

### v1

@@ I Introduction
To achieve this goal, previous research has explored human-robot collaboration in assistive scenarios  [1, 2] . Some studies  [1, 3, 4] propose modeling humans and robots as a unified system, using co-optimization to learn assistive policies where both agents are jointly trained with the same reward signal. However, this approach does not fully capture real-world dynamics. In practice, robots cannot fully access the human mind, leading to a gap or misalignment between human and robot reward functions. For the robot, the most straightforward task reward is completing the task, such as whether the food is delivered to the intended location. For humans, there is an additional layer of preference reward — preferences regarding how the task is performed, such as the speed or force used during food delivery. These human rewards are often difficult to articulate and vary across individuals, making them unknown to the robot. Therefore, robots must gather information during interactions, inferring human intentions from behavioral cues and reasoning about the underlying human preferences [5, 6, 7]. By adjusting their policies accordingly, robots can more effectively collaborate with humans to complete tasks while better aligning with human preferences and needs, thereby enhancing safety, robustness, and user satisfaction.
@@ II-A Assistive Robotics
Assistive robots have been explored as a viable solution to help disabled and elderly individuals perform daily tasks [8, 9, 10, 11, 12, 13, 1, 2, 3, 4, 14, 15]. Some works have focused on designing and developing virtual environments to facilitate the development and comparison of assistive algorithms [1, 2, 16]. For example, Assistive Gym[1] is a simulation framework for physical human-robot interaction and robotic assistance, providing several tasks crucial for daily living and introducing the co-optimization paradigm. Similarly, RCareWorld[2] proposes a framework for developing and evaluating socially-aware robotic assistance. Researchers have also designed various approaches to enhance assistive performance [17, 18, 19, 13, 20, 21, 3, 4, 22, 23]. For example, Müller et al. [17] focus on integrating human-robot collaboration to enhance ergonomics, flexibility, and quality in the assembly line, Unhelkar et al. [18] introduce a human-aware robotic assistant system that combines human motion prediction with time-optimal path planning. Gallenberger et al. [13] examine manipulation strategies for robotic feeding, focusing on how different bite acquisition and transfer methods impact the success and ease of feeding for users. Jakhotiya et al. [20] explores using recurrent neural networks to enhance policy learning for assistive tasks like feeding and dressing. Osa et al. [21] introduce a solution by maximizing mutual information between state-action pairs and latent variables, allowing for a wider range of adaptive behaviors in tasks requiring robustness. He et al. [3] aims to efficiently transfer knowledge gained from one task to another by learning robust representations, enhancing the effectiveness and versatility of assistive robots. Osa et al. [4] suggests employing a latent variable to model the behavior of an opposing agent.
@@ II-B Human-aware Policy Learning
To develop collaborative robots better suited to work alongside or for humans, some works have proposed incorporating human feelings and experiences into consideration [24, 25, 26, 27, 28, 29]. Sadigh et al. [25] propose an interactive framework that efficiently learns complex reward landscapes from minimal user input, making the learning process both efficient and aligned with human preferences while the expressiveness of such minimal user input remains limited. Chen et al. [26] present a human-in-the-loop approach for tele-operation in assistive robotics. While leveraging a control loop, their method enhances the robot’s capability to deliver real-time assistance in dynamic environments. Palan et al. [27] enables systems to understand and predict what users value without them having to specify it explicitly each time. Although these methods make valuable contributions to incorporating human factors into policy learning, they often rely on human demonstrations or preference queries, which limits their ability to autonomously learn and adapt to the complexity of human behavior.

### v2

@@ I Introduction
To achieve this goal, previous research has explored human-robot collaboration in assistive scenarios  [1, 2] . Some studies  [1, 3, 4] propose modeling humans and robots as a unified system, using co-optimization to learn assistive policies where both agents are jointly trained with the same reward signal. However, this approach does not fully capture real-world dynamics. In practice, robots cannot fully access the human mind, leading to a gap or misalignment between human and robot reward functions. For the robot, the most straightforward task reward is completing the task, such as whether the food is delivered to the intended location. For humans, there is an additional layer of preference reward — preferences regarding how the task is performed, such as the speed or force used during food delivery [1, 5]. These human rewards are often difficult to articulate and vary across individuals, making them unknown to the robot. Therefore, robots must gather information during interactions, inferring human intentions from behavioral cues and reasoning about the underlying human preferences [6, 7, 8]. By adjusting their policies accordingly, robots can more effectively collaborate with humans to complete tasks while better aligning with human preferences and needs, thereby enhancing safety, robustness, and user satisfaction.
@@ II-A Assistive Robotics
Assistive robots have been explored as a viable solution to help disabled and elderly individuals perform daily tasks [9, 10, 11, 12, 13, 14, 1, 2, 3, 4, 15, 16, 5]. Some works have focused on designing and developing virtual environments to facilitate the development and comparison of assistive algorithms [1, 2, 17]. For example, Assistive Gym[1] is a simulation framework for physical human-robot interaction and robotic assistance, providing several tasks crucial for daily living and introducing the co-optimization paradigm. Similarly, RCareWorld[2] proposes a framework for developing and evaluating socially-aware robotic assistance. Researchers have also designed various approaches to enhance assistive performance [18, 19, 20, 14, 21, 22, 3, 4, 23, 24, 5]. For example, Müller et al. [18] focus on integrating human-robot collaboration to enhance ergonomics, flexibility, and quality in the assembly line, Unhelkar et al. [19] introduce a human-aware robotic assistant system that combines human motion prediction with time-optimal path planning. Gallenberger et al. [14] examine manipulation strategies for robotic feeding, focusing on how different bite acquisition and transfer methods impact the success and ease of feeding for users. Jakhotiya et al. [21] explores using recurrent neural networks to enhance policy learning for assistive tasks like feeding and dressing. Osa et al. [22] introduce a solution by maximizing mutual information between state-action pairs and latent variables, allowing for a wider range of adaptive behaviors in tasks requiring robustness. He et al. [3] aims to efficiently transfer knowledge gained from one task to another by learning robust representations, enhancing the effectiveness and versatility of assistive robots. Osa et al. [4] suggests employing a latent variable to model the behavior of an opposing agent.
@@ II-B Human-aware Policy Learning
To develop collaborative robots better suited to work alongside or for humans, some works have proposed incorporating human feelings and experiences into consideration [25, 26, 27, 28, 29, 30, 5]. Sadigh et al. [26] propose an interactive framework that efficiently learns complex reward landscapes from minimal user input, making the learning process both efficient and aligned with human preferences while the expressiveness of such minimal user input remains limited. Chen et al. [27] present a human-in-the-loop approach for tele-operation in assistive robotics. While leveraging a control loop, their method enhances the robot’s capability to deliver real-time assistance in dynamic environments. Palan et al. [28] enables systems to understand and predict what users value without them having to specify it explicitly each time. Another line of works uses preference-based reinforcement learning (PbRL) to train agents with pairwise human feedback [31, 32, 33]. Recent advances aim to alleviate the reliance on explicit human feedback by introducing off-policy algorithms that relabel history experience and conduct unsupervised pre-training [34], and training a preference predictor to provide pseudo preference labels [35, 36, 37]. Specifically, Liu et al. [5] enhances the efficiency and robustness of preference-based reinforcement learning in complex interactive tasks by decoupling task objectives from preferences and integrating task prior knowledge into the PbRL framework. Their combination of the sketchy task reward shaping with an automated scripted teacher, reduces reliance on real human feedback during the training process. Although these methods make valuable contributions to incorporating human factors into policy learning, they still often rely on human demonstrations or explicit preference queries, which limits their ability to autonomously learn and adapt to the complexity of human behavior.